

********************************************************************
UARS WEEKLY STATUS REPORT - 22 SEP 2000
********************************************************************
***********
SOLSTICE
***********
To support spacecraft power conservation during the yaw maneuver, SOLSTICE
went into day only operations for three days last week by powering off
during orbit eclipse. Only solar observations were performed during that
time. Otherwise, SOLSTICE is in good health, collecting solar data on all
available orbital tracking periods and stellar calibration observations
whenever possible. Our solar data is not affected by the larger than normal
pointing errors (2-4 arcminutes across a track) since they are within our
calibrations.
Additionally, we continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
Level2 processing is halted due to bugs in the OA_SSPP_ATT service. CDHF
and CSC personnel are working on the problem.
***********
SUSIM
***********
SUSIM was powered down on 9/16/00 and was powered up on 9/19/00.
Because the instrument temperatures were too low, data taking did not
resume until 9/20/00. Otherwise, SUSIM continues to gather a normal
set of daily solar scans scheduled for predicted periods of UARS
telemetry.
Based on estimates of pointing accuracy during the upcoming ground
solution phase of the spacecraft attitude determination plan that will
be soon implemented, we believe that accurate solar data acquisition
by SUSIM will be problematic, i.e. uncertain. Later, when the
spacecraft-based solution becomes fully operational, we believe that
accurate solar acquisition by the SUSIM will be routine without
significant outages.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM zenith spectrometers are activate in daylight mode. PEM/AXIS
was activated at 21:15 UT on dat 264. While active, PEM/AXIS energy deposition
computation is currently suspended due to software errors. The fixes to
this code is being pursued.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
HALOE is scheduled to power on beginning-of-day September 23, 2000.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI continues to operate normally. Data collection now falls into one of
three catagories:
1) UARS low power mode. In this mode the detector is operational, but
the motors that run the filter wheels and telescope are turned off. In
this state we look straight down and measure cloud top heights. This
typically occurs for about 5 days at low beta angles.
2) HRDI full time mode. In this mode the instrument is operational for
both day and night passes and measurements are collected in the
stratosphere and mesosphere in the daytime and mesosphere at night. HRDI
is typically in this mode about a week during a yaw cycle.
3) HRDI daytime only mode. In this mode HRDI is operational during the
day when there is sufficient power. The usual science modes are run. HRDI
is in this mode when not in the other two.
Work continues on a data version 12. The changes in UARS operations and
instrument effects required significant changes in the code to account for
the dropouts. This has been completed and we are now incorporating the
effects of long term drifts into the algorithms.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG attended a meeting on Wednesday, 20 September 2000, held
to report progress on the use of Fine Sun Sensor and Three-Axis
Magnetometer data for UARS attitude control.
Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to coordinate with the Flight
Operations Team and to assist the Project Scientist in the
scheduling of instrument operations.
***********
CDHF
***********
UARS CDHF:
- The CDHF reprocessed Attitude and Orbit for UARS day 3291 and 3292
when notified level0 was ingested for Better Quality. PI was notified
for any reprocessing.
- The CDHF worked with CNE, ODIN, and NISN to debug and analyze a NISN wide
outage lasting from Friday 15 through Sunday 17 of September.
UARS CDHF REHOST:
- Successful processing on Rehost System:
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, HRDI,
MLS, SUSIM, Solstice, and WINDII.
Need Positive Response from all PI End Users prior to transition.
- SUSIM v7.2 software representative Ed Esfandiari notified
parallel testing completed. Addressing PI validation concerns of output.
- HRDI v4.4 software representative Mr. Jim Raines notified parallel
testing completed. Awaiting validation of output.
- SOLSTICE v4.62 software representative PI Barry Knapp validated
the parallel test output and stated it appears to be running properly
on the CDHFB.
- HALOE v4.82 software representative Mr. Lance Deaver notified
parallel test days completed. Awaiting validation of output.
- WINDII V5.11A software representative Mr. Brian Solheim requested parallel
processing on test days. The parallel processing is in progress.
- UARS PCG continue to monitor the rehost system for CORRELATIVE (UKMO)
and Auxiliary (SOL) data that are ingested daily.
***********
FOT
***********
This report covers 09 September 2000 (Orbit 49172, GMT Day 253) through 15
September 2000 (Orbit 49276 , GMT Day 259).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 255/0307 49203 Power Off
HALOE 257/2354 49246 Power Off
HRDI 259/2254 49275 Power Off
PEM 259/2254 49275 Power Off
The beta angle ranged between the angles of 36.1 degrees to 11.5 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
+0 degrees on 18 September 2000 (DOY 262).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.8 to 1.0 degrees C. Additional battery performance data for each
battery on 08 September 2000 and 15 September 2000 are provided in Appendix D.
The clock error ranged from -3.3 to 3.5 msec during this report period. Clock
rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 02
¥ ALERTS = 01
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 37
¥ HALOE = 01
¥ HRDI = 01
¥ PEM = 01
¥ WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 01
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/221 - 2287.496937
· Transponder B: 00/244 - 2287.499127
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/258 show the observatory data loss to be 298 hours, 16
minutes, 16.697 seconds (an increase of 2 minutes 58.840 seconds since the last
report period). This is a 0.39 percent data loss which equals a 99.61 percent
data capture for the mission.
The increase of 2 minutes 58.840 seconds occurred over five (5) days:
· 0 minutes 34.960 seconds on Day 253 due to a data hit,
· 0 minutes 25.712 seconds on Day 254 due to a data hit,
· 1 minute 28.448 seconds on Day 255 due to a late acquistion,
· 0 minutes 19.952 seconds on Day 256 due to a data hit,
· 0 minutes 9.768 seconds on Day 257 due to a data hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to operate FHST-1 for approximately one orbit each day. By
cycling the tracker on and off for such a short duration the degradation seems
to have been reduced. Scale factor calibrations are being provided on a
regular basis from FDF and are being uplinked when received. It appears that
the FOT has been able to stabilize the performance of FHST-1 for now and
continues to work on establishing another viable attitude control solution
should the tracker fail.
MEETINGS
The FOT attended a meeting to discuss FHST-1 on 11,14,15 September 2000.
The FOT attended the monthly SERB meeting on 13 September 2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for September 17, 2000 (DOY 261). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 15 Sept 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
49174 253/0443 00-472 NBTR PLAYBACK
49174 253/0529 00-472 NBTR NOOPS D-253
49181 253/1549 00-422 FHST-1 Power ON
49182 253/1648 00-472 NBTR NOOPS D-253
49183 253/1831 00-422 FHST-1 Power OFF
49183 253/1831 00-474 NBTR NOOPS D-254
49189 254/0445 00-474 NBTR PLAYBACK
49190 254/0629 00-474 NBTR NOOPS D-254
49196 254/1551 00-422 FHST-1 Power ON
49197 254/1650 00-422 FHST-1 Power OFF
49197 254/1650 00-474 NBTR NOOPS D-254
49199 254/2058 00-475 NBTR NOOPS D-255
49203 255/0307 00-476 WINDII Power OFF/Low Pwr
49204 255/0447 00-475 NBTR PLAYBACK
49205 255/0632 00-475 NBTR NOOPS D-255
49211 255/1552 00-422 FHST-1 Power ON
49212 255/1731 00-475 NBTR NOOPS D-255
49213 255/1837 00-477 NBTR NOOPS D-256
49212 255/1731 00-475 FHST-1 Power OFF
49214 255/2058 00-478 OBC GENASCII
49219 256/0449 00-477 NBTR PLAYBACK
49220 256/0635 00-477 NBTR NOOPS D-256
49226 256/1554 00-422 FHST-1 Power ON
49227 256/1654 00-477 NBTR NOOPS D-256
49228 256/1855 00-479 NBTR NOOPS D-256
49229 256/2028 00-422 FHST-1 Power OFF
49234 257/0452 00-479 NBTR Playback
49234 257/0535 00-479 NBTR NOOPS D-257
49241 257/1555 00-422 FHST-1 Power ON
49242 257/1656 00-479 NBTR NOOPS D-257
49243 257/1920 00-422 FHST-1 Power OFF
49244 257/2031 00-480 FHST-1 Scale Factor Update
49246 257/2354 00-481 HALOE Power OFF
49249 258/0454 00-482 NBTR Playback
49258 258/1923 00-483 FHST-1 Scale Factor Update/Power ON
49258 258/2034 00-422 FHST-1 Power OFF
49271 259/1558 00-422 FHST-1 Power ON
49272 259/1738 00-422 FHST-1 Power OFF
49273 259/1902 00-484 MACS Scale Factor Update
49275 259/2254 00-485 HRDI Power OFF
49275 259/2254 00-486 PEM Power OFF
_____________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 15 Sept 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIQBVOLTAGE Y-LO WINDII QB VOLT 49192
HRPSEQBIMON Y-HI PSE QB CURR MON 49192
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
OBSZ2OOL UFLTR Z2 Residual Exceeded 100 Arcsec 49229
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 15 Sept 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
-None this report period
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 15 Sept 2000
08 Aug 2000 - GMT Day 252 Beta = 36.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ----------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 14.4 0.000 0.0
2 +67.2/-16.8 +5.79/+4.51 +28.4/-15.2 24.3 1.040 21.5
3 +22.4/-56.0 +5.15/+3.87 +28.0/-16.0 24.2 1.038 21.0
15 Sept 2000 - GMT Day 259 Beta = 11.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ----------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 14.4 0.000 0.0
2 +44.8/-11.2 +5.15/+3.87 +26.0/-15.2 25.0 1.040 19.8
3 +22.4/-33.6 +4.19/+2.92 +25.6/-15.2 25.0 1.038 19.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 15 Sept 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-041 FOT 49214 255/2058 14170 Generic Late Acq.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-041 FOT see list Sept 00 14170 Generic Late Acq.
00-040 FOT 48854 231/2111 22831 Data Dropout
00-032 FHST 48211 188/2242 N/A FHST 2 Anomaly
SEVNTFW00/36